OpenStudy (anonymous):

Danika has saved enough money to pay cash for a used car. She wants to buy a car that costs $5,500 before taxes. If she will be charged 6% sales tax, what is the total amount she will pay for the car? a]$ 5,170 b]$ 5,830 c]$ 6,100

OpenStudy (anonymous):

B

OpenStudy (anonymous):

5500 * 0.06 =330 330+ 5500=5830

OpenStudy (anonymous):

.06 (decimal form of percent) multiplied by the price is your tax in dollars, add that to the price for total: .06*5500 + 5500 = 5830

OpenStudy (anonymous):

$5,500+($5,500*0.06)=$5,500+330 = $5.830